# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v1

@@ A.2 Proof of Lemma 2
$\kappa^{{\mathsf{C}}}_{2}(t)$ $\leq\log\left(1+2p^{2}(e^{t}-1)+p^{2}(e^{t}-1)^{2}\right)$
$=\log\left(1+p^{2}(e^{2t}-1)\right)\leq p^{2}(e^{2t}-1)$
@@
where $\mathrm{(a)}$ is because $x+x^{-1}\geq 2$ for any $x>0$ and $\mathrm{(b)}$ is because $\log(1+x)\leq x$ for any $x\geq 0$. Therefore, we get
$\log\mathbb{E}\left[e^{t|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|}\right]\leq\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(\gamma+2)}.$
For any $t>0$, by the Chernoff bound,
$\mathbb{P}[|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}]\leq\exp\left(-t\tau_{k}+\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(2+\gamma)}\right).$
@@
Pick $t=\frac{1}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)$. Then $t>0$ by the assumption $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$. We obtain
@@
$\mathbb{P}\left[|{\mathcal{E}}_{\pi}\cap E({\mathcal{H}}_{\pi})|\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$
A.3 Proof of Lemma 3
@@
We note that $(1+\gamma)(1-\eta)\geq 1+\gamma-\frac{\gamma}{4}>1$ and
@@
$\geq$ $(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+(1+\gamma)(-\eta)+\eta$
@@
where the last inequality is due to the fact that $(1-x)\log(1-x)+x\geq 0$ for any $10<x<$ 1 and $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. Since $\log(1+\gamma)-\frac{\gamma}{1+\gamma}\geq\frac{\gamma^{2}}{2(1+\gamma)^{2}}$, we have $\eta\gamma\leq\frac{\gamma^{2}}{4(1+\gamma)}\leq\frac{1}{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$. Therefore,
$\phi\left[(1-\eta)(1+\gamma)-1\right]\geq$ $(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]-\eta\gamma$
@@
where the last inequality is due to $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. ∎
A.4 Proof of Lemma 4
@@
For binary entropy function $h(x)=-x\log x-(1-x)\log(1-x)$, $\phi(x)=(1+x)\log(1+x)-x$ and any constant $\delta\in(0,1)$, when $m\geq\log n$,
$\sum_{k=\delta m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]\leq\left(\log n\right)^{-1+o(1)}$
@@
$\sum_{k=\delta m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]$ $\leq\sum_{k=1}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]$
$\overset{\mathrm{(a)}}{\leq}2\sum_{1\leq k\leq\frac{m}{2}}\exp\left[-k\log\left(\frac{m}{k}\right)\right]$
$\leq 2\sum_{1\leq k\leq 2\log m}\exp\left[-k\log\left(\frac{m}{k}\right)\right]+2\sum_{2\log m<k\leq\frac{m}{2}}2^{-k}$
$\leq 2\cdot\exp\left(-\log m\right)\cdot(2\log m)+2\cdot 2^{-2\log m}\overset{\mathrm{(b)}}{\leq}(\log n)^{-1+o(1)},$
where $\mathrm{(a)}$ is because $h(x)=h(1-x)$ and $h(x)\geq-x\log x$ and $\mathrm{(b)}$ is because $m\geq\log n$. ∎
A.5 Proof of Lemma 5
@@
For $P\sim\mathrm{Bern}(p,p,\rho)$ and $Q\sim\mathrm{Bern}(p,p,0)$, the KL-divergence between $P$ and $Q$ can be upper bounded by:
$D(P\|Q)\leq 25p^{2}\phi(\gamma).$
@@
By direct calculation,
$D(P\|Q)=\sum_{\{a,b\}\in\{0,1\}}p_{ab}\log\left[\frac{p_{ab}}{p^{a+b}(1-p)^{2-a-b}}\right]$
$=$ $\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\log(1-\rho)$
$+$ $\left[(1-p)^{2}+\rho p(1-p)\right]\log\left(1+\frac{\rho p}{1-p}\right)$
$\leq$ $\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\cdot(-\rho)$
$+$ $\left[(1-p)^{2}+\rho p(1-p)\right]\cdot\frac{\rho p}{1-p}$
$=$ $\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]-\rho p(1-p)+\rho^{2}\left[2p(1-p)+p^{2}\right].$
Since $\log(1+x)\geq\frac{x}{x+1}+\frac{x^{2}}{2(x+1)^{2}}$ for any $x\geq 0$, we get $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq\frac{p^{2}\gamma^{2}}{2(\gamma+1)}$. When $\gamma<3$, since $\frac{p^{2}\gamma^{2}}{2(\gamma+1)}\geq\frac{p^{2}\gamma^{2}}{8}=\frac{\rho^{2}(1-p)^{2}}{8}\geq\frac{\rho^{2}\left[2p(1-p)+p^{2}\right]}{24}$ for $0<p\leq\frac{1}{2}$, we get
$D(P\|Q)\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
When $\gamma\geq 3$, since $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq p^{2}\gamma(\log 4-1)=(\log 4-1)\rho p(1-p)$ and $\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 3\rho p(1-p)$, we get
$D(P\|Q)$ $\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]$
$\leq\left(\frac{3}{\log 4-1}+1\right)p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
Therefore, we get $D(P\|Q)\leq 25p^{2}\phi(\gamma)$. ∎
A.6 Proof of Lemma 6
@@
Suppose $\xi\sim\mathrm{Bin}(n,p)$, denote $\mu=np$, then

### v2

@@ B.2 Proof of Lemma 2
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\binom{V(G_{1})}{2}\backslash\binom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$. Applying the formulas (36) and (37) and the fact that $p_{11}\leq p$, we obtain
$\kappa^{{\mathsf{C}}}_{2}(t)$ $\leq\log\left(1+2p^{2}(e^{t}-1)+p^{2}(e^{t}-1)^{2}\right)=\log\left(1+p^{2}(e^{2t}-1)\right)\leq p^{2}(e^{2t}-1)$
@@
where $\mathrm{(a)}$ is because $p_{11}^{2}(e^{t}-1)+(e^{t}-1)^{-1}\geq 2\sqrt{p_{11}^{2}(e^{t}-1)(e^{t}-1)^{-1}}=2p_{11}$ and $\mathrm{(b)}$ is because $p_{11}=p^{2}(1+\gamma)$ and $\log(1+x)\leq x$ for any $x\geq 0$. Therefore, we obtain
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(\gamma+2)}.$
We then apply the Chernoff bound the provide an upper bound for $\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]$. For any $t>0$,
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-t\tau_{k}+\frac{|{\mathcal{E}}_{\pi}|}{2}p^{2}(e^{2t}-1)+\frac{k\gamma}{4(2+\gamma)}\right).$
@@
Let $t=\frac{1}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)$. Then $t>0$ by the assumption $\tau_{k}>|\mathcal{E}_{\pi}|p^{2}$. We obtain that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}+\frac{k\gamma}{4(2+\gamma)}\right).$
B.3 Proof of Lemma 3
We first upper bound the higher-order cumulants. Let $t=\frac{\rho}{3\sqrt{1+\rho^{2}}}$. Then $0<t<\frac{1}{1+\rho}$ since $10<\rho<$ 1. By Lemma 1,
@@ B.3 Proof of Lemma 3
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})}\right]\leq\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\binom{V(G_{1})}{2}\backslash\binom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$. Applying the formulas (41) and (42), we obtain that
$\kappa_{1}^{\mathsf{C}}(t)-\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)=$ $\frac{1}{4}\log\left[\frac{\left(1-t^{2}(1-\rho^{2})-2t\rho\right)\left(1-t^{2}(1-\rho^{2})+2t\rho\right)}{\left(1-t^{2}(1-\rho^{2})-2t\rho\right)^{2}}\right]$
$=$ $\frac{1}{4}\log\left[1+\frac{4t\rho}{1-t^{2}(1-\rho^{2})-2t\rho}\right]\overset{\mathrm{(a)}}{\leq}\frac{1}{4}\log\left(1+\frac{12\rho^{2}}{9-7\rho^{2}+\rho^{4}}\right)\overset{\mathrm{(b)}}{\leq}\frac{\log 5}{4},$
where $\mathrm{(a)}$ is because $\frac{4t\rho}{1-t^{2}(1-\rho^{2})-2t\rho}=\frac{12\rho^{2}}{9\sqrt{1+\rho^{2}}-\frac{\rho^{2}(1-\rho^{2})}{\sqrt{1+\rho^{2}}}-6\rho^{2}}\leq\frac{12\rho^{2}}{9-\rho^{2}(1-\rho^{2})-6\rho^{2}}=\frac{12\rho^{2}}{9-7\rho^{2}+\rho^{4}}$; $\mathrm{(b)}$ is because $\frac{12x^{2}}{9-7x^{2}+x^{4}}$ is increasing on $(0,1)$ and $\rho<1$.
By the Chernoff bound, we have that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right)\right).$
It remains to upper bound $-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
$-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$ $=-t\tau_{k}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left[1-\frac{2\rho^{2}}{9}+\frac{\rho^{4}(1-\rho^{2})^{2}}{81(1+\rho^{2})^{2}}\right]$
$\overset{\mathrm{(a)}}{\leq}-t\tau_{k}+\frac{9|\mathcal{E}_{\pi}|}{28}\left[\frac{2\rho^{2}}{9}-\frac{\rho^{4}(1-\rho^{2})^{2}}{81(1+\rho^{2})^{2}}\right]\overset{\mathrm{(b)}}{\leq}-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14},$
where $\mathrm{(a)}$ is because $\frac{2\rho^{2}}{9}-\frac{\rho^{4}(1-\rho^{2})^{2}}{81(1+\rho^{2})^{2}}=\frac{\rho^{2}\left(18(1+\rho^{2})^{2}-\rho^{2}\left(1-\rho^{2}\right)^{2}\right)}{81(1+\rho^{2})^{2}}\geq 0$ and $\log(1-x)\geq-\frac{9x}{7}$ for any $0\leq x\leq\frac{2}{9}$; $\mathrm{(b)}$ is because $t\geq\frac{\rho}{6}$. Therefore, we obtain that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{6}+\frac{\rho^{2}|\mathcal{E}_{\pi}|}{14}+\frac{\log 5}{8}k\right).$
B.4 Proof of Lemma 4
We first upper bound the higher-order cumulants. By Lemma 1, for any $t>0$,
$\log\mathbb{E}\left[e^{t\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)}\right]\leq\frac{|{\mathcal{E}}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right),$
where $L$ denotes the number of self-loops. The self-loop for $e$ only happens when $\pi(e)=\pi^{*}(e)$. For $uv\in\binom{V(G_{1})}{2}\backslash\binom{F_{\pi}}{2}$, by the definition of $F_{\pi}$, $\pi(u)\neq\pi^{*}(u)$ or $\pi(v)\neq\pi^{*}(v)$. Therefore, $\pi(uv)=\pi^{*}(uv)$ implies that $\pi(u)=\pi^{*}(v)$ and $\pi(v)=\pi^{*}(u)$. Since $d(\pi^{*},\pi)=k$, we must have $L\leq\frac{k}{2}$. Let $t=\frac{\rho}{4(1-\rho)}$. Applying the formulas (45) and (46), we obtain that
$\kappa_{1}^{\mathsf{C}}(t)-\frac{1}{2}\kappa_{2}^{\mathsf{C}}(t)$ $=\frac{1}{4}\log\left[\frac{(1+2t-2t\rho)(1+2t+2t\rho)}{(1+2t(1-\rho))^{2}}\right]$
$=\frac{1}{4}\log\left(1+\frac{4t\rho}{1+2t-2t\rho}\right)\leq\frac{1}{4}\log\left(\frac{1}{1-\rho}\right),$
where the last inequality follows from $1+\frac{4t\rho}{1+2t-2t\rho}=\frac{2+\rho^{2}-\rho}{2+\rho}\cdot\frac{1}{1-\rho}\leq\frac{1}{1-\rho}$.
By the Chernoff bound, we have that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}\left(\mathcal{H}_{\pi}^{f}\right)\geq\tau_{k}\right]\leq\exp\left(-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa^{{\mathsf{C}}}_{2}(t)+L\left(\kappa^{{\mathsf{C}}}_{1}(t)-\frac{1}{2}\kappa^{{\mathsf{C}}}_{2}(t)\right)\right).$
It remains to upper bound $-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$. We note that
$-t\tau_{k}+\frac{|\mathcal{E}_{\pi}|}{2}\kappa_{2}^{\mathsf{C}}(t)$ $=-t\tau_{k}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left[(1+2t)^{2}-4t^{2}\rho^{2}\right]$
$\leq-t\tau_{k}-\frac{|\mathcal{E}_{\pi}|}{4}\log(1+4t)=-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right).$
Therefore, we obtain that
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right).$
Appendix C Auxiliary results
@@
We note that $(1+\gamma)(1-\eta)\geq 1+\gamma-\frac{\gamma}{4}>1$, and
@@
$\geq(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+(1+\gamma)(-\eta)+\eta$
@@
where the last inequality is due to the fact that $(1-x)\log(1-x)+x\geq 0$ for any $10<x<$ 1 and $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. Since $\eta\leq\frac{\gamma}{4(1+\gamma)}$ and $(1+\gamma)\log(1+\gamma)-\gamma\geq\frac{\gamma^{2}}{2(1+\gamma)}$, we obtain $\eta\gamma\leq\frac{\gamma^{2}}{4(1+\gamma)}\leq\frac{1}{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$. Therefore,
$\phi\left[(1-\eta)(1+\gamma)-1\right]$ $\geq(1-\eta)\left[(1+\gamma)\log(1+\gamma)-\gamma\right]-\eta\gamma$
@@
where the last inequality is because $0<\eta\leq\frac{\gamma}{4(1+\gamma)}<\frac{1}{4}$. ∎
@@
For any $m\geq 10$,
$\sum_{k=1}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]\leq\frac{4\log m+2}{m},$
where $h(x)=-x\log x-(1-x)\log(1-x)$ is the binary entropy function.
@@
$\sum_{k=1}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]$ $\overset{\mathrm{(a)}}{\leq}2\sum_{1\leq k\leq\frac{m}{2}}\exp\left[-k\log\left(\frac{m}{k}\right)\right]$
$\overset{\mathrm{(b)}}{\leq}2\sum_{1\leq k\leq 2\log m}\exp\left[-k\log\left(\frac{m}{k}\right)\right]+2\sum_{2\log m+1\leq k\leq\frac{m}{2}}2^{-k}$
$\overset{\mathrm{(c)}}{\leq}2\cdot\exp\left(-\log m\right)\cdot(2\log m)+2\cdot 2^{-2\log m}$
$\overset{\mathrm{(d)}}{\leq}\frac{4\log m+2}{m},$
where $\mathrm{(a)}$ is because $h(x)=h(1-x)$ and $h(x)\geq-x\log x$; $\mathrm{(b)}$ is because $\log\left(\frac{m}{k}\right)\geq\log 2$ when $k\leq\frac{m}{2}$; $\mathrm{(c)}$ is because $k\log\left(\frac{m}{k}\right)\geq\log m$ for $1\leq k\leq 2\log m$ when $m\geq 10$; $\mathrm{(d)}$ is because $2\cdot 2^{-2\log m}\leq\frac{2}{m}$. ∎
@@
For $P_{1}=\mathrm{Bern}(p,p,\rho)$ and $Q_{1}=\mathrm{Bern}(p,p,0)$, we have
$D(P_{1}\|Q_{1})\leq 25p^{2}\phi(\gamma).$
For $P_{2}=\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ and $Q_{2}=\mathcal{N}\left(\begin{bmatrix}0\\ 0\end{bmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)$, we have
$D(P_{2}\|Q_{2})=\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$
@@
Recall $p_{ab}$ for any $a,b\in\{0,1\}$ defined in (4). Then, we have
$D(P_{1}\|Q_{1})=$$~{}\sum_{\{a,b\}\in\{0,1\}}p_{ab}\log\left[\frac{p_{ab}}{p^{a+b}(1-p)^{2-a-b}}\right]$
$=$ $~{}\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\log(1-\rho)$
$~{}+\left[(1-p)^{2}+\rho p(1-p)\right]\log\left(1+\frac{\rho p}{1-p}\right)$
$\overset{\mathrm{(a)}}{\leq}$ $~{}\left[p^{2}+\rho p(1-p)\right]\log\left[1+\frac{\rho(1-p)}{p}\right]+2p(1-p)(1-\rho)\cdot(-\rho)$
$~{}+\left[(1-p)^{2}+\rho p(1-p)\right]\cdot\frac{\rho p}{1-p}$
=$\overset{\mathrm{(b)}}{=}$ $~{}p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right],$
where $\mathrm{(a)}$ uses $\log(1+x)\leq x$ for any $x>-1$; $\mathrm{(b)}$ follows from $\gamma=\frac{\rho(1-p)}{p}$. Since $\log(1+x)\geq\frac{x}{x+1}+\frac{x^{2}}{2(x+1)^{2}}$ for any $x\geq 0$, we obtain that $p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]\geq\frac{p^{2}\gamma^{2}}{2(\gamma+1)}$. When $\gamma<3$, since $\frac{p^{2}\gamma^{2}}{2(\gamma+1)}\geq\frac{p^{2}\gamma^{2}}{8}=\frac{\rho^{2}\cdot 3(1-p)^{2}}{24}\geq\frac{\rho^{2}\left[2p(1-p)+p^{2}\right]}{24}$ for $0<p\leq\frac{1}{2}$, we obtain that
$D(P_{1}\|Q_{1})\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
When $\gamma\geq 3$, since $\rho^{2}\left[2p(1-p)+p^{2}\right]\leq 3\rho p(1-p)$ and $(\log 4-1)\rho p(1-p)=p^{2}\gamma(\log 4-1)\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$, we obtain
$D(P_{1}\|Q_{1})$ $\leq p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]+\rho^{2}\left[2p(1-p)+p^{2}\right]$
$\leq\left(\frac{3}{\log 4-1}+1\right)p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right]$
$\leq 25p^{2}\left[(1+\gamma)\log(1+\gamma)-\gamma\right].$
Therefore, we get $D(P_{1}\|Q_{1})\leq 25p^{2}\phi(\gamma)$.
We denote $P_{2}(a,b)$ and $Q_{2}(a,b)$ the probability density function under $P_{2}$ and $Q_{2}$, respectively. Then, the KL-divergence between $P_{2}$ and $Q_{2}$ is given by
$D(P_{2}\|Q_{2})$ $=\iint P_{2}(a,b)\log\left(\frac{P_{2}(a,b)}{Q_{2}(a,b)}\right)\,\mathrm{d}a\mathrm{d}b$
$=\iint P_{2}(a,b)\left[\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right)+\frac{\rho ab}{1-\rho^{2}}-\frac{\rho^{2}(a^{2}+b^{2})}{2(1-\rho^{2})}\right]\,\mathrm{d}a\mathrm{d}b$
$=\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right)+\frac{\rho^{2}}{1-\rho^{2}}-\frac{2\rho^{2}}{2(1-\rho^{2})}=\frac{1}{2}\log\left(\frac{1}{1-\rho^{2}}\right).\qed$
@@
Suppose $\xi\sim\mathrm{Bin}(n,p)$, denote $\mu=np$, for any $\delta>0$,
@@
Lemma 9 (Hanson-Wright inequality).
Let $X,Y\in\mathbb{R}^{n}$ be standard Gaussian vectors such that the pairs $(X_{i},Y_{i})\sim\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ are independent for $i=1,\cdots,n$. Let $M_{0}\in\mathbb{R}^{n\times n}$ be any deterministic matrix. There exists some universal constant $c_{0}>0$ such that,
$\mathbb{P}\left[\left|X^{\top}M_{0}Y-\rho\mathrm{Tr}(M_{0})\right|\geq c_{0}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]\leq\delta.$
Proof.
Note that $X^{\top}M_{0}Y=\frac{1}{4}(X+Y)^{\top}M_{0}(X+Y)-\frac{1}{4}(X-Y)^{\top}M_{0}(X-Y)$ and
$\mathbb{E}\left[(X+Y)^{\top}M_{0}(X+Y)\right]=(2+2\rho)\mathrm{Tr}(M_{0}),\mathbb{E}\left[(X-Y)^{\top}M_{0}(X-Y)\right]=(2-2\rho)\mathrm{Tr}(M_{0}).$
By Hanson-Wright inequality [HW71], there exists some universal constant $c_{0}$ such that
$\mathbb{P}\left[\left|\frac{1}{4}(X+Y)^{\top}M_{0}(X+Y)-\frac{2+2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]\leq\frac{\delta}{2},$
$\mathbb{P}\left[\left|\frac{1}{4}(X-Y)^{\top}M_{0}(X-Y)-\frac{2-2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]\leq\frac{\delta}{2}$
for any $\delta>0$. Consequently,
$~{}\mathbb{P}\left[\left|X^{\top}M_{0}Y-\rho\mathrm{Tr}(M_{0})\right|\geq c_{0}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]$
$\leq$ $~{}\mathbb{P}\left[\left|\frac{1}{4}(X+Y)^{\top}M_{0}(X+Y)-\frac{2+2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]$
$+\mathbb{P}\left[\left|\frac{1}{4}(X-Y)^{\top}M_{0}(X-Y)-\frac{2-2\rho}{4}\mathrm{Tr}(M_{0})\right|\geq\frac{c_{0}}{2}\left(\|M_{0}\|_{F}\sqrt{\log(1/\delta)}\vee\|M_{0}\|_{2}\log(1/\delta)\right)\right]$
$\leq$ $~{}\delta.\qed$
Lemma 10 (Chernoff’s inequality for Chi-squared distribution).
Suppose $\xi$ follows the chi-squared distribution with $n$ degrees of freedom. Then, for any $\delta>0$,
$\mathbb{P}\left[\xi>(1+\delta)n\right]\leq\exp\left(-\frac{n}{2}\left(\delta-\log\left(1+\delta\right)\right)\right).$ (53)
Proof.
The result follows from [Gho21, Theorem 1]. ∎